# Search is All You Need for Few-shot Anomaly Detection

### v1

@@ 3.4. Category-Indexed Memory Bank
During the testing phase, the category of the test image is first retrieved by measuring the similarity between the class token $\mathbf{G}_{s}$ of the test image and the stored tokens. The corresponding patch features can then be retrieved through the category, enabling anomaly detection and localization, as shown in Figure 2. The pseudo-code for establishing both the patch feature memory bank and the global feature memory bank is provided in Algorithm 2.
@@ 4.3. Visualization Results
Fig. 5 shows the partial visualization results of VisionAD on MVTec and VisA dataset. Compared to PatchCore (Roth et al., 2022) and WinCLIP (Jeong et al., 2023), VisionAD produces more distinctive and precise anomaly maps in the 1-shot setting. Additionally, VisionAD is capable of accurately locating subtle anomalies that are often overlooked. This robustness and reliability under few-shot settings underscore the potential practical applicability of VisionAD in real-world scenarios. For more qualitative results refer to Appendix C.
@@ Appendix C Qualitative Visualization
Appendix D Failure Cases Study
To ensure a fair comparison with the INP on the Real-IAD dataset, we similarly selected only 1, 2, or 4 images per category as support samples. However, during the analysis of anomaly location (visualizations), we observed that VisionAD produced accurate results for images of the sample captured from similar views, whereas its performance noticeably degraded when the views varied significantly, as shown in Figure LABEL:realiad_failur. This observation confirms our initial hypothesis that the notion of 1-/2-/4-shot should fundamentally refer to the number of samples rather than the number of images. Therefore, future research should consider collecting at least five distinct views per sample to construct the support set.
@@ Appendix D Failure Cases Study
Table A2. Scaling input size on MVTec-AD (%). †: default.

### v2

@@ 3.4. Category-Indexed Memory Bank
During the testing phase, the category of the test image is first retrieved by measuring the similarity between the class token $\mathbf{G}_{s}$ of the test image and the stored tokens. The corresponding patch features can then be retrieved through the category, enabling anomaly detection and localization, as shown in Figure 2.
@@ 4.3. Visualization Results
Fig. 5 shows the partial visualization results of VisionAD on MVTec and VisA dataset. Compared to PatchCore (Roth et al., 2022) and PromptAD (Li et al., 2024), VisionAD produces more distinctive and precise anomaly maps in the 1-shot setting. Additionally, VisionAD is capable of accurately locating subtle anomalies that are often overlooked. This robustness and reliability under few-shot settings underscore the potential practical applicability of VisionAD in real-world scenarios. For more qualitative results refer to Appendix C.
@@ Appendix C Qualitative Visualization
Appendix D Additional Dataset
To further assess the applicability and generalization of the proposed CRR, we apply it to two real-world industrial tasks under both class-separated and multi-class settings, including the CSDD and HSS-IAD datasets. The CSDD (Wang et al., 2024a) (Casting Surface Defect Detection) dataset aims to bridge the gap between real-world defect detection and the idealized conditions of existing datasets. It contains 12,647 high-resolution grayscale images with pixel-precise ground truth annotations for all defect samples. The HSS-IAD (Heterogeneous Same-Sort Industrial Anomaly Detection) dataset includes 8,580 images of metallic-like industrial parts with detailed anomaly annotations. The parts in this dataset exhibit diverse structures and appearances, with subtle defects that closely resemble the base materials.
In Table A3, the quantitative results on CSDD and HSS-IAD under different settings are presented. As the number of support images increases, anomaly detection and segmentation performance significantly improve. Under the multi-class setting, VisionAD struggles to accurately distinguish among casting categories of HSS-IAD, as the casting surfaces share highly similar semantics. This leads to a slight drop in the mean performance on HSS-IAD. For CSDD, it is also challenging to accurately find the corresponding support set for each test image due to the semantic similarity among samples of the same class. Therefore, we do not report performance results for this setting. Considering the complexity and high accuracy required in real-world scenarios, the performance of current FSAD methods still falls short of meeting practical industrial demands.
Appendix E Failure Cases Study
To ensure a fair comparison with the INP on the Real-IAD dataset, we similarly selected only 1, 2, or 4 images per category as support samples. However, during the analysis of anomaly location (visualizations), we observed that VisionAD produced accurate results for images of the sample captured from similar views, whereas its performance noticeably degraded when the views varied significantly, as shown in Figure A1. This observation confirms our initial hypothesis that the notion of 1-/2-/4-shot should fundamentally refer to the number of samples rather than the number of images. Therefore, future research should consider collecting at least five distinct views per sample to construct the support set.
@@ Appendix E Failure Cases Study
Table A2. Scaling input size on MVTec-AD and VisA (%). †: default.
@@
Table A3. Performance on CSDD and HSS-IAD under multi-class and class-separated UAD setting (%). CS: Class-Separated Setting. MC: Multi-Class Setting.
Setting Setup CSDD HSS-IAD
AUROC AUPR pAUROC PRO AUROC AUPR pAUROC PRO
CS 1-shot 48.5$\pm$1.8 57.1$\pm$1.2 69.9$\pm$1.8 27.5$\pm$1.3 68.8$\pm$1.1 74.1$\pm$1.1 85.4$\pm$0.6 65.8$\pm$1.1
2-shot 49.8$\pm$2.2 57.6$\pm$1.9 69.9$\pm$1.9 27.3$\pm$2.1 69.2$\pm$0.9 74.0$\pm$0.6 85.7$\pm$0.2 65.6$\pm$0.8
4-shot 50.3$\pm$0.9 57.4$\pm$0.9 73.0$\pm$0.8 30.9$\pm$2.3 69.6$\pm$1.1 73.7$\pm$1.3 86.7$\pm$0.3 66.9$\pm$0.5
8-shot 53.0$\pm$1.7 58.9$\pm$1.2 75.3$\pm$0.8 34.8$\pm$2.1 71.7$\pm$1.7 74.8$\pm$1.2 87.5$\pm$0.5 67.6$\pm$1.2
MC 1-shot - - - - 67.7$\pm$1.0 73.7$\pm$1.1 84.2$\pm$1.7 63.3$\pm$2.9
2-shot - - - - 67.8$\pm$1.4 73.2$\pm$1.3 84.6$\pm$1.4 63.2$\pm$1.9
4-shot - - - - 67.8$\pm$1.8 72.8$\pm$1.4 85.3$\pm$0.7 64.5$\pm$1.0
8-shot - - - - 69.0$\pm$2.0 73.3$\pm$1.4 85.8$\pm$0.9 64.7$\pm$2.0